How many 3 letter alphanumeric passwords can you generate if you can't begin or end the password with the letter O, but any digit can repeat within the password? Also, the password cannot be ABC?
Answer #1

For 1st and 3rd digit, we have 25 choices and for 2nd, we have 26 choices. So,

Total number of choices available = 25 * 26 * 25 = 16250

In this, there would be one passsword as ABC so total number of possible combinations = 16250 - 1 = 16249